OpenStudy (anonymous):

How many meat balls are required for a party of 20 adults and 5 children the formula is m=4a+2c , where c=number of children; m= number of meatballs; a= number of adults

OpenStudy (anonymous):

give me a second

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay so you have to plug in the number of adults(a) and the number of children(c) to the equation. There are 20 adults(a) and 5 children(c). So then we can solve by multiplying 4 by 20 and 2 by 5. Then we will have m=90. Does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay good! (www.instagram.com/coryshearer_official)

OpenStudy (texaschic101):

use your formula : m = 4a + 2c adults (a) = 20 children (c) = 5 now just sub m = 4(20) + 2(5) m = 80 + 10 m = 90 Its that simple :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank yall

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

how can it be 90

OpenStudy (texaschic101):

a = number of adults and there are 20 c = number of children and there are 5 formula : m = 4a + 2c apparently the formula is set up where every adult gets 4 meatballs and every child gets 2 meatballs. so we just sub m = 4a + 2c m = 4(20) + 2(5) m = 80 + 10 m = 90

OpenStudy (texaschic101):

does that help ?